# Linear Complexity Computation of Code Distance and Minimum Size of Trapping Sets for LDPC Codes with Bounded Treewidth

### v1

@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
$g_{t}(R,Q)=\begin{cases}g_{t^{\prime}}(R,Q),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\}\text{and }R=\Gamma_{o}(G_{t}[v]);\\ g_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\}),&\text{otherwise.}\end{cases}$ (4)
First, we prove that the left-hand side of (3) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$.
If $(R,Q)$ satisfies $v\notin Q$, then $v\notin S$ and $S\subset\mathcal{L}(G_{t^{\prime}})$ is also a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap(B_{t}^{v}\setminus\{v\})\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S]),S\cap B_{t}^{v}\Big{)}\\ &=(R,Q)\subset\left(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}\right).\end{split}$ (5)
This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
If $(R,Q)$ satisfies $Q=\{v\}$ and $R=\Gamma_{o}(G_{t}[v])$, then $S$ is either $\{v\}$ or $\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set with $b=0$ in $G_{t}^{\prime}$ such that $S_{t^{\prime}}\cap B_{t^{\prime}}=\emptyset$. In this case, we have $f_{t}(R,Q)=1$ and $g_{t}(R,Q)=1$.
For other cases of $(R,Q)$, $S_{t}$ has the form $S=\{v\}\cup S_{t^{\prime}}$, where $S_{t^{\prime}}$ is a trapping set in $G_{t^{\prime}}$. This implies that the parameters of $S_{t^{\prime}}$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S_{t^{\prime}}]),S_{t^{\prime}}\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S\setminus\{v\}]),(S\setminus\{v\})\cap(B_{t}^{v}\setminus\{v\})\Big{)}\\ &\overset{Prop.1}{=}\Big{(}\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G_{t}[v]),(S\cap B_{t}^{v})\setminus(\{v\}\cap\{v\})\Big{)}\\ &\overset{(b)}{=}\Big{(}R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\}\Big{)}\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}).\end{split}$ (6)
The reason why $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$ holds in (b) is that $v$ belongs only to $B_{t}$. According to the definition of tree decomposition, $v$ can only be adjacent to vertices in $B_{t}^{c}$ within $G_{t}$. This leads to $\mathcal{N}_{G_{t}}(v)\subset B_{t}^{c}$, which results $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$. Therefore, $|S|=|S_{t^{\prime}}|+1$ implies $f_{t}(R,Q)\geq f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\})+1$.
On the other hand, we now proceed to prove that the left-hand side of (3) is less than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, if $(R,Q)$ satisfies $v\notin Q$, let $R^{\prime}=R\subset B_{t^{\prime}}^{c}$ and $Q^{\prime}=Q\subset B_{t^{\prime}}^{v}$, and $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$. Then, $S_{t^{\prime}}$ remains a trapping set in $G_{t}$. By reversing the steps in (5), we obtain that $S_{t^{\prime}}$ retains the parameters $(R^{\prime},Q^{\prime})$ in $G_{t}$, which leads to $f_{t}(R^{\prime},Q^{\prime})\leq f_{t^{\prime}}(R^{\prime},Q^{\prime})$. If $(R,Q)$ satisfies $Q=\{v\}$ and $R\neq\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$ or $v\in Q\neq\{v\}$, let $R^{\prime}=R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])\subset B_{t^{\prime}}^{c},Q^{\prime}=Q\backslash\{v\})\subset B_{t^{\prime}}^{v}$, $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$, and $S=\{v\}\cup S_{t^{\prime}}$ be a trapping set in $G_{t}$. By applying the reverse derivation of (6), we obtain that the parameters of $S$ in $G_{t}$ remain as $(R^{\prime}\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q^{\prime}\cup\{v\})$. This leads to $f_{t^{\prime}}(R^{\prime},Q^{\prime})+1\geq f_{t}(R^{\prime}\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q^{\prime}\cup\{v\}).$ Thus, we have $f_{t}(R,Q)\leq f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v\cup B_{t}^{c}]),Q\setminus\{v\})+1.$
It takes $O(2^{k})$ time for each introduced variable node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Forget variable node: Suppose $t$ is a forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{v\}$, where $v$ is a variable node. This means the functions $f_{t}$ and $g_{t}$ are restrictions of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively. Specifically, $f_{t}$ and $g_{t}$ are defined over a smaller domain $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}\setminus\{v\}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\},$ (7)
@@
$\begin{split}g_{t}(R,Q)&=\mathbb{1}_{f_{t}(R,Q)=f_{t^{\prime}}(R,Q)}g_{t}(R,Q)\\ &+\mathbb{1}_{f_{t}(R,Q)=f_{t^{\prime}}(R,Q\cup\{v\})}g_{t}(R,Q\cup\{v\}),\end{split}$ (8)
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where $\mathbb{1}_{x=y}$ is the indicator function,
$\mathbb{1}_{x=y}=\begin{cases}1,&\text{if $x=y$};\\ 0,&\text{if $x\neq$y}.\end{cases}$ (9)
Next, we provide proof for the validity of Equation (7). For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S],S\cap B_{t^{\prime}}^{v}\Big{)}\\ &=\Big{(}G_{t^{\prime}}[S],S\cap(B_{t}^{v}\cup\{v\})\Big{)}\\ &\overset{Prop.4}{=}\Big{(}G_{t}[S],(S\cap B_{t}^{v})\cup(S\cap\{v\})\Big{)}\\ &=\Big{(}R,Q\cup(S\cap\{v\})\Big{)}.\end{split}$ (10)
If $v\notin S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. If $v\in S$, then the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q\cup\{v\})\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. Therefore,
$f_{t}(R,Q)\geq\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\}.$
Conversely, for any trapping set $S_{t^{\prime}}$ in $G_{t^{\prime}}$ with parameters $(R,Q)$ or $(R,Q\cup\{v\})$, its parameters in $G_{t}$ can be derived as $(R,Q)$ by applying Equation (10). This implies that
$f_{t}(R,Q)\leq\min\{f_{t^{\prime}}(R,Q),f_{t^{\prime}}(R,Q\cup\{v\})\}.$
It takes $O(2^{k})$ time for each forget variable node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
introduced check node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{c\}$, where $c$ is a check node. This means the functions $f_{t}$ and $g_{t}$ are extensions of the domains of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively, as $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}\cup\{c\}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\begin{cases}f_{t^{\prime}}(R\setminus\{c\},Q),&\text{if }c\in R\text{, $E(c,Q)$ is odd;}\\ f_{t^{\prime}}(R,Q),&\text{if }c\notin R\text{, $E(c,Q)$ is even;}\\ +\infty,&\text{otherwise},\end{cases}$ (11)
@@
$g_{t}(R,Q)=\begin{cases}g_{t^{\prime}}(R\setminus\{c\},Q),&\text{if }c\in R\text{, $E(c,Q)$ is odd;}\\ g_{t^{\prime}}(R,Q),&\text{if }c\notin R\text{, $E(c,Q)$ is even;}\\ 0,&\text{otherwise},\end{cases}$ (12)
for all $R\in 2^{B_{t}^{c}}$, $Q\in 2^{B_{t}^{v}}$, and $d\in\{0,1,...,b\}$, where $E(c,Q)$ denotes the number of edges between $c$ and $Q$ in $G_{t}$, and it is easy to compute as it only requires determining the degree of $c$ in $G_{t}[Q\cup\{c\}]$. Specifically, $E(c,Q)$ being even includes the case where $E(c,Q)=0$.
First, we prove that the left-hand side of (11) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. Since $\mathcal{L}(G_{t})=\mathcal{L}(G_{t^{\prime}})$, $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c}\setminus\{c\})]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.2}{=}\Big{(}\Gamma_{o}\big{(}G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c})]\big{)}\oplus\Gamma_{o}\big{(}G_{t}[S\cup\{c\}]\big{)},Q\Big{)}\\ &\overset{(c)}{=}\Big{(}\Gamma_{o}(G_{t}[S])\oplus\Gamma_{o}(G[Q\cup\{c\}]),Q\Big{)}\\ &=\Big{(}R\oplus\Gamma_{o}(G[Q\cup\{c\}]),Q\Big{)}.\end{split}$ (13)
The validity of equation (c) follows from the definition of tree decomposition, which implies that $c$ can only be adjacent to variable nodes in $B_{t}^{v}$ and not to those in $S\setminus B_{t}^{v}$. Consequently, the number of edges between $c$ and $S$ is equal to the number of edges between $c$ and $S\cap B_{t}^{v}=Q$.
If $c$ and $S$ have an odd number of edges in $G_{t}$, then $c\in R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\{c\}$. In this case, the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R\setminus\{c\},Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R\setminus\{c\},Q)$.
If $c$ and $S$ have an even number of edges in $G_{t}$, then $c\notin R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\emptyset$. Consequently, the parameters of $S$ in $G_{t^{\prime}}$ are given by $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. This leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
There are no other possible cases of trapping sets in $G_{t}$. Hence, $f_{t}(R,Q)=\infty$ for any other cases of $R$ and $Q$.
As for the proof that the left-hand side of (11) is less than or equal to the right-hand side, consider each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$. If $(R,Q)$ satisfies $c\in R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\{c\}$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R\setminus\{c\},Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. By (13), $S_{t^{\prime}}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Therefore, $f_{t}(R,Q)\leq f_{t^{\prime}}(R\setminus\{c\},Q)$. If $(R,Q)$ satisfies $c\notin R$ and $\Gamma_{o}(G[Q\cup\{c\}])=\emptyset$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v})$. By (13), $S_{t^{\prime}}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Therefore, $f_{t}(R,Q)\leq f_{t^{\prime}}(R,Q)$.
It takes $O(2^{k})$ time for each introduced check node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Forget check node: Suppose $t$ is an forget node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\setminus\{c\}$, where $c$ is a check node. This means the functions $f_{t}$ and $g_{t}$ are restrictions of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively. Specifically, $f_{t}$ and $g_{t}$ are defined over a smaller domain $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}\setminus\{c\}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=f_{t}^{\prime}(R,Q),$ (14)

### v2

@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
$g_{t_{i}}(L,I)=\begin{cases}g_{t_{j}}(L,I),&\text{if }v\notin L;\\ 1,&\text{if }L=\{v\},\text{ and }I=\Gamma_{G_{i}}^{o}(v);\\ g_{t_{j}}(L\backslash\{v\},I\oplus\Gamma_{G_{i}}^{o}(v)),&\text{otherwise.}\end{cases}$ (2)
Next, we prove the correctness of (1) and (2). Let $S$ be a trapping set in $G$. For each pair $(L,I)$, the following conclusion holds.
  • If $v\notin L$, $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S$ is also a $(L,I,t_{j})$-type trapping set, which leads to $f_{t_{i}}(L,I)=f_{t_{j}}(L,I)$ and $g_{t_{i}}(L,I)=g_{t_{j}}(L,I)$. This is because, when $v\notin S$, the following equation holds:
    $\begin{split}S\subset\mathcal{L}_{j}\Leftrightarrow S\subset\mathcal{L}_{i}\end{split}$ (3)
    $\begin{split}S\cap B_{t_{j}}&=S\cap(B_{t_{i}}\setminus\{v\})\\ &=(S\cap B_{t_{i}})\setminus(S\cap\{v\})\\ &=S\cap B_{t_{i}}\end{split}$ (4)
@@
    $\Gamma_{G_{j}}^{o}(S)=\Gamma_{G_{i}}^{o}(S).$ (5)
    When $S$ is a $(L,I,t_{i})$-type trapping set, we have $S\cap B_{t_{i}}=L$, and $v\notin L$, thus $v\notin S$. When $S$ is a $(L,I,t_{j})$-type trapping set, we have $S\subset\mathcal{L}_{j}$, thus $v\notin S$.
  • If $L=\{v\}$ and $I=\Gamma^{o}_{G_{i}}(v)$, then $\{v\}$ is a $(L,I,t_{i})$-type trapping set. In this case, we have $f_{t_{i}}(L,I)=1$ and $g_{t_{i}}(L,I)=1$.
  • For other cases, $S$ has the form $S=\{v\}\cup S^{\prime}$, where $v\notin S^{\prime}$. We claim $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S^{\prime}$ is a $(L\setminus\{v\},I\oplus\Gamma^{o}_{G_{i}}(v),t_{j})$-type trapping set, which leads to $f_{t_{i}}(L,I)=f_{t_{j}}(L\setminus\{v\},I\oplus\Gamma^{o}_{G_{i}}(v))+1.$ and $g_{t_{i}}(L,I)=g_{t_{j}}(L\setminus\{v\},I\oplus\Gamma^{o}_{G_{i}}(v))$. According to $B_{t_{i}}=B_{t_{j}}\cup\{v\}$, we have:
    $\begin{split}S^{\prime}\subset\mathcal{L}_{j}\Leftrightarrow S\subset\mathcal{L}_{i}\end{split}$ (6)
    $\begin{split}S^{\prime}\cap B_{t_{j}}&=(S\setminus\{v\})\cap(B_{t_{i}}\setminus\{v\})\\ &=(S\cap B_{t_{i}})\setminus(\{v\}\cap\{v\})\\ &=(S\cap B_{t_{i}})\setminus\{v\}\end{split}$ (7)
@@
    $\begin{split}\Gamma^{o}_{G_{j}}(S^{\prime})&=\Gamma^{o}_{G_{i}}(S\setminus\{v\})\\ &\overset{prop.1}{=}\Gamma^{o}_{G_{i}}(S)\oplus\Gamma^{o}_{G_{i}}(v)\end{split}$ (8)
    Clearly, based on (6) $\sim$ (8), the sufficiency of the claim is established. Conversely, to establish necessity, suppose $S^{\prime}$ is a $(L\setminus{v},I\oplus\Gamma^{o}_{G_{i}}(v),t_{j})$-type trapping set, then $S\cap B_{t_{i}}=L$ or $S\cap B_{t_{i}}=L\setminus\{v\}$. Since $v\in S$, it follows that $S\cap B_{t_{i}}=L$. Therefore, $S$ is a $(L,I,t_{i})$-type trapping set.
It takes $O(2^{k})$ time for each introduced variable node to compute $f_{t_{i}}(L,I)$ and $g_{t_{i}}(L,I)$ for all $(L,I)$.
Forget variable node: Suppose $t_{i}$ is a forget node with a unique child $t_{j}$, and that the corresponding bags satisfy $B_{t_{i}}=B_{t_{j}}\setminus\{v\}$, where $v$ is a variable node. We can compute $f_{t_{i}}$ and $g_{t_{i}}$ using the following equations for all $L\subset{\mathcal{L}^{B}_{i}}$, $I\subset{\mathcal{R}^{B}_{i}}$:
$f_{t_{i}}(L,I)=\min\{f_{t_{j}}(L,I),f_{t_{j}}(L\cup\{v\},I)\}.$ (9)